OpenStudy (anonymous):

Determine whether the following function is even, odd, or neither. Please show work. f(x) = x^2/3 + x -1 Determine whether the following function is even, odd, or neither. Please show work. f(x) = x^2/3 + x -1 @Mathematics

OpenStudy (anonymous):

a function is even when f(x) =f(-x) it is odd when f(-x) = -f(x) in other case neither odd nor even X = 8 as a example now f(-8) = 4 - 8 -1= -5 f(8) = 4+8-1 = 11 so it neigher satisfies any above conditions hence it is neither odd nor even

OpenStudy (anonymous):

Did you mean to say neither? You wrote "neigher", I assume a typo?

OpenStudy (anonymous):

ya it is neither.. !!

OpenStudy (anonymous):

ok- so my answer to this question would be neither?

OpenStudy (anonymous):

yes!!, your answer is f(x) is neither odd nor even function

OpenStudy (anonymous):

ok great. Sorry- this stuff confuses me so much. Thank you for helping!

OpenStudy (anonymous):

welcome!!